# MIDI: Multi-Instance Diffusion for Single Image to 3D Scene Generation

### v1

@@ 1 Introduction
Recent image-to-3D object generation models [27, 37, 40, 41, 23, 60, 25, 71, 65, 33, 64, 68, 61, 63, 36, 73, 66], with powerful 3d prior and generalization capabilities, can generate high-quality geometry from a single object image. Building upon these pre-trained models, a common approach for scene generation involves using them as tools within a multi-step compositional generation process, which includes segmenting the scene image, completing individual object images, generating each object, and optimizing their spatial relationships [74, 5, 20], as shown in Fig. 2. While these methods leverage the priors of 3D object generation models, the generation process is inherently lengthy and prone to error accumulation – errors in intermediate steps can significantly distort the final result. Moreover, the optimization of spatial relationships cannot directly optimize 3D objects generated one by one by the previous stage that lacks global scene context, leading to misalignments between the generated instances and the overall scene. Therefore, if inter-object spatial relationships can be modeled directly within the 3D generation model, it is possible to construct an end-to-end pipeline that addresses these issues by generating all instances simultaneously with coherent spatial arrangements.
@@ 2.1 Scene Reconstruction from a Single Image
Recovering the 3D structure of a scene from a single image is a fundamental challenge in computer vision. Existing methods can be categorized into feed-forward reconstruction methods [45, 70, 47, 7, 35, 18, 72, 6, 4], retrieval-based methods [19, 26, 30, 31, 32, 17], and recent compositional generation approaches [74, 5, 20, 11].
@@
Advancements in diffusion models [22, 57] and large-scale datasets [9, 10] have propelled progress in 3D generation [37, 40, 41, 23, 60, 25, 71, 65, 33, 64, 68, 61, 63, 36, 66, 73, 53, 67, 43, 34, 12]. Several image-to-3D object generation methods [40, 41, 60, 64, 68, 63, 61] adopt a two-stage pipeline that involves generating multi-view images and then reconstructing 3D objects. They fine-tune pre-trained image [54, 49] or video [2] generation models to produce multi-view images and employ feed-forward reconstruction models [23, 60, 69, 75] or optimization-based methods [62] to recover geometries. Another group of work [71, 33, 66, 73] focuses on generating 3D native geometry by training large-scale generative models, which typically comprise a variational autoencoder [28] and a latent diffusion transformer (DiT) [48]. These models produce high-quality geometries with strong generalization due to training on diverse datasets. Building upon these advancements, we fine-tunes such an object geometry generator to create compositional instances while retaining generalization ability.
@@ 3 Preliminary: 3D Object Generation Models
Large-scale 3D object generation models [73, 71, 66, 33] are the foundation of our approach. These models often comprise three main components: 1) a transformer-based variational autoencoder (VAE) [28] with an encoder $\mathcal{E}$ and a decoder $\mathcal{D}$, which compress 3D geometric representations into a low-dimensional latent space, and 2) a denoising transformer network $\epsilon_{\theta}$, trained on the compressed latent space to transform noise $\bm{\epsilon}\sim\mathcal{N}(0,I)$ into the original 3D data distribution $\mathbf{z}_{0}$, and 3) a group of condition encoders, such as CLIP [50] and DINO encoders [46] for encoding text or image conditions, which are then passed to the denoising network by cross-attention mechanism.
@@ 4.3 Training
To train MIDI, we extend the loss of our base model, which utilizes the rectified flow [39] architecture, from single-object to multi-instance. For each training step, we sample a shared noise level $t$ from $1$ to $T$ for all the instances $\{\mathbf{z}^{i}\}_{i=1}^{N}$, perturbing them along a simple linear trajectory:
@@ 6 Conclusion
Future works.  Building upon our proposed multi-instance diffusion that utilizes pre-trained object generation models for compositional 3D scene generation, future work can explore several directions: 1) extending the approach to model more complex interactions in compositional scenes, such as characters interacting with objects (e.g. “a panda playing a guitar”), which requires specialized datasets; 2) incorporating explicit 3D geometric knowledge to develop more efficient and expressive multi-instance attention mechanisms; 3) investigating the latent, implicit 3D perception capabilities of scene generation models; and 4) scaling the framework to handle a larger number of objects and operate in open-world environments.
@@
Supplementary Material
Figure 9: Detailed comparison between existing compositional generation methods and our multi-instance diffusion.
7 Background
Base model.  Following scalable 3D object generation methods [73, 71, 66, 33], we firstly trains a VAE to compress 3D geometric representations into a low-dimensional latent space. Specifically, $\mathbf{x}\in\mathbb{R}^{L\times 6}$, which represents positions and normals of $L$ points, are mapped to latent space by $\mathbf{z}=\mathcal{E}(\mathbf{x})$, where $\mathbf{z}\in\mathbb{R}^{l\times c}$, and $l$ denotes the length of the tokens after compression. The latents are converted back to the 3D space by regressing signed distance function (SDF) values using $\mathbf{s}=\mathcal{D}(\mathbf{z})$.
Next, the denoising network $\epsilon_{\theta}$ is trained in the compressed latent space to transform noise $\bm{\epsilon}\sim\mathcal{N}(0,I)$ into the original 3D data distribution. During training, following the rectified flow architecture [39], the original data $\mathbf{z}_{0}$ is perturbed along a simple linear trajectory:
$\mathbf{z}_{t}=t\mathbf{z}_{0}+(1-t)\bm{\epsilon}$ (4)
for $t=1,\cdots,T$, where $T$ represents the number of steps in the diffusion process. In practice, we adopt logit-normal sampling [14] to increase the weight for intermediate steps. The denoising network $\epsilon_{\theta}$, featuring 21 attention blocks with residual connections, is trained to approximate the slope of the distribution transformation trajectory by minimizing the following loss:
$\mathbb{E}_{\mathbf{z},\mathbf{y},\bm{\epsilon}\sim\mathcal{N}(0,I),t}[\lVert\mathbf{z}_{0}-\bm{\epsilon}-\epsilon_{\theta}(\mathbf{z}_{t},t,\tau_{\theta}(\mathbf{y}))\rVert_{2}^{2}]$ (5)
where $\tau_{\theta}$ is the image encoder, and $\mathbf{y}$ is the conditioning image, incorporated into the denoising transformer via cross-attention mechanism.
8 Implementation Details
Training.  we trained MIDI to simultaneously generate up to $N=5$ instances. We selected this value based on an analysis of the 3D-FRONT dataset [15], where we observed that scenes containing five or fewer objects constitute the majority, while scenes with more than five objects are relatively rare. Instead of excluding scenes with more than $5$ objects, we employed a clustering method to select five representative objects from such scenes for training. During training, we randomly dropped the image conditioning with a probability of 0.1. We adopted the same strategy as in the training of the base model, utilizing logit-normal sampling [14] to increase the weight of intermediate diffusion steps, which helps the model focus on the more challenging stages of the generation process. For the training configuration, we used a learning rate of $5\times 10^{-5}$ and trained MIDI for 5 epochs on 8 NVIDIA A100 GPUs.
Inference.  In our experimental setup, we first used Grounded-SAM [52] to segment the scene images, obtaining masks for individual objects. We then applied our multi-instance diffusion model to generate compositional 3D instances using classifier-free guidance [21], which enhances the fidelity and coherence of the generated scenes. We set the number of inference steps to 50 and the guidance scale to 7.0. The entire process of generating a 3D scene from a single image takes approximately 40 seconds on an NVIDIA A100 GPU.
9 Additional Discussions
MIDI vs. compositional generation methods.  As show in Fig. 9, existing compositional generation methods involve a multi-step process, generating 3D objects one by one and then optimizing their spatial relationships. However, this type of methods lack the contextual information of the global scene when generating objects, thus generating inaccurate or mismatched 3D objects. In addition, it is very difficult to optimize the accurate scene layout based on a single image, and the position of similar objects will be reversed when there are similar objects in the scene (as shown in Fig. 9). In contrast, our method models object completion, 3D generation and spatial relationships in a multi-instance diffusion model, thus generating coherent and accurate 3D scenes.
10 Limitations
While MIDI generates 3D instances within the global scene coordinate system—specifically, a normalized space ranging from $-1$ to $1$—this approach causes smaller objects to occupy a relatively minor portion of the overall space. Consequently, these small objects may have lower resolution compared to objects generated in their canonical spaces, where the entire capacity of the model can focus on a single object. We believe that enhancing the multi-instance diffusion model to generate objects in their canonical spaces, along with their spatial positions within the scene, could address this issue by allowing each object to be generated at optimal resolution.
Also, our model is constrained by the simplicity of interaction relationships present in existing scene datasets. As a result, MIDI may struggle to generate scenes featuring intricate interactions, such as objects with dynamic interplays. We anticipate that introducing more complex and diverse training data, encompassing a wider variety of object interactions and spatial relationships, would enhance the model’s capacity to generalize at the level of object spatial interactions. This improvement would enable the generation of scenes with more sophisticated and realistic inter-object dynamics.

### v2

@@ 1 Introduction
Recent image-to-3D object generation models [27, 38, 41, 42, 22, 62, 25, 75, 68, 33, 67, 71, 64, 66, 37, 77, 69, 63, 19, 73], with powerful 3d prior and generalization capabilities, can generate high-quality geometry from a single object image. Building upon these pre-trained models, a common approach for scene generation involves using them as tools within a multi-step compositional generation process, which includes segmenting the scene image, completing individual object images, generating each object, and optimizing their spatial relationships [78, 5, 20], as shown in Fig. 2. While these methods leverage the priors of 3D object generation models, the generation process is inherently lengthy and prone to error accumulation – errors in intermediate steps can significantly distort the final result. Moreover, the optimization of spatial relationships cannot directly optimize 3D objects generated one by one by the previous stage that lacks global scene context, leading to misalignments between the generated instances and the overall scene. Therefore, if inter-object spatial relationships can be modeled directly within the 3D generation model, it is possible to construct an end-to-end pipeline that addresses these issues by generating all instances simultaneously with coherent spatial arrangements.
@@ 2.1 Scene Reconstruction from a Single Image
Recovering the 3D structure of a scene from a single image is a fundamental challenge in computer vision. Existing methods can be categorized into feed-forward reconstruction methods [46, 74, 48, 7, 36, 17, 76, 6, 4], retrieval-based methods [18, 26, 30, 31, 32, 16], and recent compositional generation approaches [78, 5, 20, 11, 61].
@@
Advancements in diffusion models [21, 58] and large-scale datasets [9, 10] have propelled progress in 3D generation [38, 41, 42, 22, 62, 25, 75, 68, 33, 67, 71, 64, 66, 37, 69, 77, 54, 70, 44, 35, 12]. Several image-to-3D object generation methods [41, 42, 62, 67, 71, 66, 64, 24] adopt a two-stage pipeline that involves generating multi-view images and then reconstructing 3D objects. They fine-tune pre-trained image [55, 50] or video [2] diffusion models to produce multi-view images and employ large reconstruction models [22, 62, 72, 79] or optimization-based methods [65] to recover geometries. Another group of work [75, 33, 69, 77, 34] focuses on generating 3D native geometry by training large-scale generative models, which typically comprise a variational autoencoder [28] and a latent diffusion transformer (DiT) [49]. These models produce high-quality geometries with strong generalization due to training on diverse datasets. Building upon these advancements, we fine-tunes such an object geometry generator to create compositional instances while retaining generalization ability.
@@ 3 Preliminary: 3D Object Generation Models
Large-scale 3D object generation models [77, 75, 69, 33, 34] are the foundation of our approach. These models often comprise three main components: 1) a transformer-based variational autoencoder (VAE) [28] with an encoder $\mathcal{E}$ and a decoder $\mathcal{D}$, which compress 3D geometric representations into a low-dimensional latent space, and 2) a denoising transformer network $\epsilon_{\theta}$, trained on the compressed latent space to transform noise $\bm{\epsilon}\sim\mathcal{N}(0,I)$ into the original 3D data distribution $\mathbf{z}_{0}$, and 3) a group of condition encoders, such as CLIP [51] and DINO encoders [47] for encoding text or image conditions, which are then passed to the denoising network by cross-attention mechanism.
@@ 4.3 Training
To train MIDI, we extend the loss of our base model, which utilizes the rectified flow [40] architecture, from single-object to multi-instance. For each training step, we sample a shared noise level $t$ from 00 to $1$ for all the instances $\{\mathbf{z}^{i}\}_{i=1}^{N}$, perturbing them along a simple linear trajectory:
@@ 6 Conclusion
Limitations and future works.  MIDI performs relatively poorly for tiny-resolution image input and complex inter-instance interaction, as shown in the supplementary materials. Building upon our proposed multi-instance diffusion for compositional 3D scene generation, future work can explore several directions: 1) extending the approach to model more complex interactions in compositional scenes, such as characters interacting with objects (e.g. “a panda playing a guitar”), which requires specialized datasets; 2) incorporating explicit 3D geometric knowledge to develop more efficient and expressive multi-instance attention mechanisms; 3) investigating the latent, implicit 3D perception capabilities of scene generation models; and 4) scaling the framework to handle a larger number of objects and operate in open-world environments.
@@
Acknowledgment
This work was supported by National Natural Science Foundation of China (62132001), and the Fundamental Research Funds for the Central Universities.